OpenStudy (anonymous):

medal Simplify. 14 – 6.2 ∙ (–3 + –7) A. –78 B. –48 C. 76 D. 78

OpenStudy (i_need_help_please):

14−6.2(−3−7) =14−(6.2)(−10) =14−(−62) =76

OpenStudy (anonymous):

thank you!!!

OpenStudy (i_need_help_please):

You're Welcome